OpenStudy (anonymous):

Amy has a triangular flag. She wants to cut out a similar flag that has dimensions one-sixth of the original flag. If the area of the original flag is 2160 square inches, what is the area of the new flag that Amy cuts out? 60 in2 50 in2 20 in2 10 in2

OpenStudy (anonymous):

60 in^2

OpenStudy (anonymous):

60

OpenStudy (anonymous):

The two flags will be similar. If two polygons are similar, the ratio of their areas is equal to the square of the scale factor of the two polygons. (1/6)^2 = A/2160 where A is the area of the smaller flag 1/36 = A/2160 36A = 2160 A = 60 square inches